OpenStudy (anonymous):

Is the following problem simplified correctly? (3x+1)(2x-3)=6x^2-9x-3

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\( -3-7 x+6 x^2 \)

OpenStudy (anonymous):

6x^2-7x-3

OpenStudy (anonymous):

,. use foil method to make it eaier to solve. multiply the First terms(of binomial terms). (3x)*(2x) = 6x^2..right'? ---this will be the first term of your simplified answer. then the Outer term(of binomial terms)... (3x)*(-3) = -9x.. nxt is the Inner term(of binomial terms). (1)(2x) = 2x.. but dont forget you must add the Outer term result with the Inner term result... -9x + 2x = -7x( the second term of your simplified answer. lastly You. have to multiply the Lasts terms(of binomial terms)... 1*-3 = -3.. Add all the results. then you will arrive at the answer.. 6x^2 - 7x -3